# A General Framework to Evaluate Methods for Assessing Dimensions of Lexical Semantic Change Using LLM-Generated Synthetic Data

### v1

@@ Appendix G Multilevel Modeling Approach
To determine the most appropriate model, we compare the null model, simplified random intercepts model, and random slopes model using the Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC), which provide measures of model fit, with lower values indicating better balance between fit and complexity. Likelihood ratio tests assess whether including additional random effects significantly improve model fit. Higher Log Likelihood (LL) indicates better fit.
Model Diagnostics:
Residual diagnostics are performed on the final model to ensure key assumptions are met:
  • Normality: Q-Q plots; Shapiro-Wilk test.
  • Homoscedasticity: Residual vs. fitted value plots; Levene’s test for homogeneity of variances.
  • Random Effects Variance: Variance estimates for random intercepts and residuals to quantify group-level variability contribution.
Model results are summarized in Table 12, showing that increasing levels of synthetic injections significantly increase Dimension indices.
Method IL $\mathbf{\beta}$ $\mathbf{SE}$ $\mathbf{p}$ LL $\sigma^{2}$
Valence Index - -0.001 0.000 <.0001 59.85 0.02
+ 0.003 0.000 <.0001 56.30 0.03
Cosine Distance - NA NA NA NA NA
+ <0.0001 <0.0001 <.0001 94.36 0.001
Arousal Index - -0.002 0.000 <.0001 63.45 0.01
+ 0.002 0.000 <.0001 68.38 0.009
@@
Table 12: Results of the Final Mixed Linear Models Predicting Dimension Scores from Injection Levels
Note: IL = Injection level. $\beta$ = Regression coefficient for synthetic injection level. SE = standard errors. p-values test the null hypothesis that the coefficient is zero. LL = Log-Likelihood, indicating model fit. $\sigma^{2}$ = Variance (Random Effects), quantifies variability due to grouping. NA = Not available.
Sentiment (Positive)
  • The Null Model showed an ICC of 0.59, indicating that 59% of variance in avg_valence_index_positive is attributable to target, justifying its inclusion as a random intercept.
[…]
  • Based on model comparison (Log-Likelihood: Simplified Model = 56.30, Random Slopes Model = 45.40; AIC/BIC unavailable due to convergence issues), the Simplified Model was selected as the final model (see Table 13).
Measure Value
Number of Observations (Groups) 36 (6)
Log-Likelihood 56.30
Scale 0.0006
Random Effects Variance (Intercepts) 0.026
Fixed Effect (injection_level) 0.003
   SE $\pm$ 0.000
   z 27.49
   p-value < 0.0001
Table 13: Model with Random Intercepts Predicting Valence Index from Positive Sentiment Injection Level.
Sentiment (Negative)
  • The Null Model showed an ICC of 0.884, indicating that 88.4% of variance in avg_valence_index_negative is attributable to target, justifying its inclusion as a random intercept.
[…]
  • Given model comparison results (LL: Simplified Model = 59.85, Random Slopes Model = 48.90; AIC/BIC unavailable due to convergence issues), the Simplified Model was selected as the final model (see Table 14).
Measure Value
Number of Observations (Groups) 36 (6)
Log-Likelihood 59.85
Scale 0.0005
Random Effects Variance (Intercepts) 0.021
Fixed Effect (injection_level) -0.001
   SE $\pm$ 0.000
   z -11.40
   p-value < 0.0001
Table 14: Model with Random Intercepts Predicting Valence Index from Negative Sentiment Injection Level.
Intensity (High)
  • The analyses assess the impact of injection_level on avg_arousal_index_high. The Null Model (ICC = 0.54) justified including target as a random intercept, as 54% of variance was attributable to group differences.
[…]
  • A Random Slopes Model, allowing injection_level to vary across groups, failed to converge due to overparameterization or insufficient data. Its exceptionally low scale (0.0002) suggested overfitting or model specification issues.
  • Based on convergence and parsimony, the Simplified Model with Random Intercepts was selected as the final model. LL metrics (Simplified: 68.38, Random Slopes: 59.99) confirmed this choice (see Table 15).
Measure Value
Number of Observations (Groups) 36 (6)
Log-Likelihood 68.38
Scale 0.0003
Random Effects Variance (Intercepts) 0.009
Fixed Effect (injection_level) 0.002
   SE $\pm$ 0.000
   z 23.63
   p-value <.0001
Table 15: Model with Random Intercepts Predicting Arousal Index from High Intensity Injection Level.
Intensity (Low)
  • The Null Model (ICC = 0.53) justified including target as a random intercept, as 53% of variance was attributable to group differences.
[…]
  • A Random Slopes Model, allowing injection_level to vary across groups, failed to converge due to overparameterization or insufficient data. Its exceptionally low scale (0.0002) suggested overfitting or model specification issues.
  • The Simplified Model with Random Intercepts was selected as the final model, supported by LL metrics (Simplified: 63.45, Random Slopes: 58.63). See Table 16.
Measure Value
Number of Observations (Groups) 36 (6)
Log-Likelihood 63.45
Scale 0.0004
Random Effects Variance (Intercepts) 0.011
Fixed Effect (injection_level) -0.002
   SE $\pm$ 0.000
   z -22.98
   p-value <.0001
Table 16: Model with Random Intercepts Predicting Arousal Index from Low Intensity Injection Level.
Breadth
  • The analyses assess the impact of injection_level on cosine_distance_mean. The Null Model (ICC = 0.71) justified including target as a random intercept, as 71% of variance was attributable to group differences.
[…]
  • A Random Slopes Model, allowing injection_level to vary across groups, failed to converge, likely due to overparameterization or insufficient data. The model’s scale was near zero (<0.0001), suggesting overfitting or misspecification.
  • Based on convergence and parsimony, the Simplified Model with Random Intercepts was selected as the final model. LL metrics (Simplified: 94.36, Random Slopes: 84.01) confirmed this choice (see Table 17).
Measure Value
Number of Observations (Groups) 36 (6)
Log-Likelihood 94.36
Scale 0.0001
Random Effects Variance (Intercepts) 0.001
Fixed Effect (injection_level) <0.001
   SE $\pm$ <0.0001
   z 7.49
   p-value <.0001
Table 17: Model with Random Intercepts Predicting Arousal Index from High Intensity Injection Level.
Appendix H SIB Scores: Results for Five-Year Random Sampling Strategy
@@
Appendix I Alternative LSC Detection Methods: Results for Bootstrapped Settings

### v2

@@ Appendix H Linear Mixed-Effects Modeling for SIB Dimensions
Likelihood ratio tests (LRTs) favored the random slopes model across dimensions, but due to convergence issues and minimal slope variance, we selected the more stable and parsimonious random intercepts model for interpretation.
Model Results Summary:
We fit separate mixed-effects models for each outcome (Valence, Arousal, Breadth), with injection_level as a fixed effect and random intercepts for target. Outcomes were standardized (z-scores) to ensure coefficient comparability. Model assumptions (linearity, homoscedasticity, normality) were verified via standard diagnostics. Table 12 reports standardized coefficients ($\beta$), 95% confidence intervals, and p-values. All models showed significant, directionally consistent effects. Random intercept variance ($\sigma^{2}$) indicates the extent of baseline score differences across target terms.
Score IL $\mathbf{\beta\ (95\%\ CI)}$ $\mathbf{p}$ $\sigma^{2}$
Valence + 0.611   (0.567, 0.654) <.0001 0.732
- -0.305   (-0.357, -0.253) <.0001 1.059
Arousal + 0.638   (0.585, 0.691) <.0001 0.681
- -0.643   (-0.698, -0.588) <.0001 0.672
Breadth + 0.429   (0.317, 0.541) <.0001 0.845
@@
Table 12: Results of the Final Mixed Linear Models Predicting Dimension Scores from Injection Levels.
Note: IL = Injection level (– = increased variation, + = decreased variation). $\beta$ = Standardized coefficient with 95% confidence interval. p-values test the null hypothesis that the coefficient is zero. $\sigma^{2}$ = Random intercept variance. Number of Observations (Groups) 36 (6).
Appendix I SIB Scores: Results for Five-Year Random Sampling Strategy
@@
Appendix J Alternative LSC Detection Methods: Results for Bootstrapped Settings